Pyridine is a weak base that is used in the manufacture of pesticides and plastic resins. It is also a component of cigarette smoke. Pyridine ionizes in water as follows: C5H5N+H2O⇌C5H5NH++OH− The pKb of pyridine is 8.75. What is the pH of a 0.200 M solution of pyridine? (Assume that the temperature is 25 ∘C.) Express the pH numerically to two decimal places.
Answer #1

The pKb of pyridine is 8.75

PKb = 8.75

-logKb = 8.75

Kb   = 1.78*10^-9

------- C5H5N+H2O⇌C5H5NH++OH−

I ------ 0.2 -----------   0 ------- 0

C----- -x    ------------- +x ------- +x

E --- 0.2-x ------------ +x ------ +x

          Kb    = [C5H5NH^+][OH^-]/[C5H5N]

         1.78*10^-9   = x*x/(0.2-x)

         1.78*10^-9*(0.2-x) = x^2

         x   = 1.89*10^-5

    [OH^-]   = x   = 1.89*10^-5M

    POH = -log[OH^-]

              = -log(1.89*10^-5)

              = 4.72

PH      = 14-POH

            = 14-4.72

            = 9.28 >>>.answer
Answer #2

Your answer is actually the pKa, in order to find the pH you need to find the -log[H^+].

source: Mastering Chemistry
answered by: berrygrace